# Advancing GUI for Generative AI: Charting the Design Space of Human-AI Interactions through Task Creativity and Complexity

### 04 Jan 2024 v1

@@ Abstract.
conference: IUI’24: 29th International Conference on Intelligent User Interfaces; March 18 – 21, 2024; Greenville, South Carolinabooktitle: IUI’24: 29th International Conference on Intelligent User Interfaces, March 18 – 21, 2024, Greenville, South Carolinajournalyear: 2023copyright: rightsretainedconference: 28th International Conference on Intelligent User Interfaces; March 18–21, 2024; Greenville, South Carolina, United Statesbooktitle: 28th International Conference on Intelligent User Interfaces (IUI ’24 Companion), March 18–21, 2024, Greenville, South Carolina, United Statesdoi: 10.1145/3581754.3584111isbn: 979-8-4007-0107-8/23/03
@@ 2.1. Fixed-scope Content Curation Tasks
Substantial research demonstrates that Large Language Models (LLMs) adeptly manage defined content curation tasks such as text summarization (Dang et al., 2022; Goyal et al., 2022), content refinement (Lin, 2023), and code explanation (MacNeil et al., 2022). These tasks involve reiterating existing knowledge succinctly and coherently without generating novel insights. GPT-3 and similar LLMs have shown proficiency in these tasks, often without the necessity for human intervention (Brown et al., 2020). Clark et al. revealed that text generated by GPT-3 was linguistically advanced to the extent that distinguishing it from human-written text became challenging for evaluators (Clark et al., 2021). This calls for well-structured frameworks to dissect human and machine errors and ascertain authorship (Dou et al., 2022). Such evidence underscores the potential for reducing human involvement in content curation tasks due to the superior quality of machine-generated text.
@@ 2.2. Atomic Creative Tasks
The second category comprises atomic creative tasks that necessitate generating novel and valuable outputs (Sawyer, 2012; Runco and Jaeger, 2012). These include tasks such as crafting analogies, metaphors, analogous design concepts (Zhu and Luo, 2022; Bhavya et al., 2022; Lee et al., 2022), slogans (Clark et al., 2018), and tweetorials (Gero et al., 2022). LLMs like GPT-3 can foster novel connections or ”creative leaps” due to their extensive knowledge base (Chan et al., 2017). However, real creativity often requires domain-specific and nuanced knowledge that might be absent in LLM training data. Hence, to ensure quality, LLM outputs for these tasks must be guided, selected, or edited by humans through precise human-AI interactions.
@@ 2.3. Complex and Interdependent Tasks
Beyond the atomic creative tasks are complex and interdependent tasks such as text-to-image generation (Brade et al., 2023; Chung and Adar, 2023), and multifaceted tasks like storytelling in text (Yuan et al., 2022; Singh et al., 2022) and images (Yan et al., 2023). These tasks demand not only domain expertise but also capabilities for planning, reasoning, ideating, and maintaining context over extended periods to create coherent and innovative content. Researchers have developed tools like Promptify (Brade et al., 2023) and PromptPaint (Chung and Adar, 2023) for text-to-image generation, Spellburst (Angert et al., 2023) for exploratory creative coding, and XCreation (Yan et al., 2023) for cross-modal storytelling to facilitate these intricate interactions.
@@ 3.1. Study 1 on Fixed-Scope Content Curation Tasks: News Headline Generation
This research investigates human-AI interactions within fixed-scope content curation tasks, specifically focusing on the generation of news headlines (Ding et al., 2023a). A between-subjects experiment was conducted involving 40 participants tasked with creating news headlines using either traditional manual methods or one of three LLM-enhanced systems: (1) selection, (2) guidance + selection, and (3) guidance + selection + post-editing:
@@ 4. Implications and future work
Our impending research aims to explore further into the intricacies of human-AI interactions, particularly focusing on complex and interdependent tasks that exhibit a high potential for interaction. This study builds upon the insights gained from Study 3, which primarily addressed the early stages of data visualization prototyping facilitated by generative AI. This process often resulted in a disorganized, ”messy” canvas, as evidenced by the results of Study 3. We now shift our focus to the subsequent phase: effectively presenting data analysis findings from these visualizations to the end-readers.
To enhance the interpretability of the data visualization results, we will use Generative AI to transform the non-linear, dispersed results on the canvas into a linear, easy-to-read narrative consisting of text and visualizations. This transformation will be based on: 1) the visualizations and their associated information, including prompts used to generate those visualizations, descriptions (captions) of insights from those visualizations, the order in which these visualizations were created, and their positions on the canvas; 2) information about the readers, encompassing their interests, backgrounds, and prior knowledge. Similar to Study 1, we will conduct a between-subjects study to test various human-AI interactions aimed at linearizing visualizations. These interactions include:
  • AI only: AI converts the canvas to a linear narrative using only the authoring information.
  • Selection: AI generates multiple narratives with authoring information, which the author of the visualizations then selects.
  • Post-Editing: AI generates one narrative with authoring information, which is then post-edited by the author of the visualizations.
  • Author Guidance: AI generates one narrative with both authoring information and keywords provided by the author of the visualizations.
  • End-Reader Guidance: AI generates one narrative with authoring information and incorporates the preferences, interests, and prior knowledge of the end-readers of the visualizations.
  • Author Only: The author of the visualizations crafts the narrative without AI assistance.
Through this comprehensive evaluation, we aim to develop a more structured and intuitive method for presenting complex data analysis results from the exploration phase, making them accessible and understandable to a wider audience.

### 12 Feb 2024 v3

@@ Abstract.
conference: IUI’24: 29th International Conference on Intelligent User Interfaces; March 18 – 21, 2024; Greenville, South Carolinabooktitle: IUI’24: 29th International Conference on Intelligent User Interfaces, March 18 – 21, 2024, Greenville, South Carolinajournalyear: 2024copyright: rightsretainedconference: 29th International Conference on Intelligent User Interfaces - Companion; March 18–21, 2024; Greenville, SC, USAbooktitle: 29th International Conference on Intelligent User Interfaces - Companion (IUI Companion ’24), March 18–21, 2024, Greenville, SC, USAdoi: 10.1145/3640544.3645241isbn: 979-8-4007-0509-0/24/03
@@ 2.1. Fixed-scope Content Curation Tasks
Substantial research demonstrates that Large Language Models (LLMs) adeptly manage defined content curation tasks such as text summarization (Goyal et al., 2022), content refinement (Lin, 2023), and code explanation (MacNeil et al., 2022). These tasks involve reiterating existing knowledge succinctly and coherently without generating novel insights. GPT and similar LLMs have shown proficiency in these tasks, often without the necessity for human intervention. Clark et al. revealed that text generated by GPT-3 was linguistically advanced to the extent that distinguishing it from human-written text became challenging for evaluators (Clark et al., 2021). This calls for well-structured frameworks to dissect human and machine errors and ascertain authorship (Dou et al., 2022). Such evidence underscores the potential for reducing human involvement in content curation tasks due to the superior quality of machine-generated text.
@@ 2.2. Atomic Creative Tasks
The second category comprises atomic creative tasks that necessitate generating novel and valuable outputs (Sawyer, 2012; Chan et al., 2024). These include tasks such as crafting analogous design concepts (Zhu and Luo, 2022; Lee et al., 2022), design problems (MacNeil et al., 2021b, a, c), slogans (Clark et al., 2018), and tweetorials (Gero et al., 2022). LLMs like GPT can foster novel connections or ”creative leaps” due to their extensive knowledge base (Chan et al., 2017). However, real creativity often requires domain-specific and nuanced knowledge that might be absent in LLM training data. Hence, to ensure quality, LLM outputs for these tasks must be guided, selected, or edited by humans through precise human-AI interactions.
@@ 2.3. Complex and Interdependent Tasks
Beyond the atomic creative tasks are complex and interdependent tasks such as active search (Palani et al., 2021b, a), neurocognitive disorder test (Ding et al., 2022), collaborative design work (MacNeil et al., 2023), and storytelling in text (Yuan et al., 2022; Singh et al., 2022) and images (Yan et al., 2023). These tasks demand not only domain expertise but also capabilities for planning, reasoning, ideating, and maintaining context over extended periods to create coherent and innovative content. Researchers have developed tools like Promptify (Brade et al., 2023) and PromptPaint (Chung and Adar, 2023) for text-to-image generation, Spellburst (Angert et al., 2023) for exploratory creative coding, and XCreation (Yan et al., 2023) for cross-modal storytelling to facilitate these intricate interactions.
@@ 3.1. Study 1 on Fixed-Scope Content Curation Tasks: News Headline Generation
This research investigates human-AI interactions within fixed-scope content curation tasks, specifically focusing on the generation of news headlines (Ding et al., 2023a). A between-subjects experiment was conducted involving 40 participants tasked with creating news headlines using either traditional manual methods or one of three LLM-enhanced systems: (1) selection, (2) guidance + selection, and (3) guidance + selection + post-editing.
@@ 4. Implications and future work
Our upcoming research aims to explore further into the intricacies of human-AI interactions, particularly focusing on complex and interdependent tasks that exhibit a high potential for interaction. This research builds upon the findings from Study 3, which investigated the preliminary stages of data visualization prototyping facilitated by Generative AI. These initial attempts often led to an unstructured, ”messy” output, as demonstrated by the outcomes of Study 3. Moreover, our objective extends beyond achieving a neatly organized set of data visualizations. Data analysis aims to uncover new insights, such as explanations for discrepancies observed when comparing visualizations. While the interpretation of these visualizations traditionally falls within the expertise of human data analysts, the advent of generative AI and its multi-modal understanding capabilities, including image captioning, presents an opportunity for AI to aid in interpreting and structuring visualizations. Thus, our research question is: How can we harness multi-modal understanding of generative AI for real-time data visualization management? To answer this question, we plan to conduct a formative study focusing on how experienced data analysts interpret and organize data visualizations as Knudsen et al. (Knudsen et al., 2012). This formative study will be followed by the development of a mixed-initiative user interface designed for real-time data visualization management, and a user study to evaluate the performance of the system.